OpenStudy (anonymous):

help me graph this?? 3x + 5y < 10 x – 4y 12

OpenStudy (anonymous):

its suposed to be two lines that intercept each other

OpenStudy (anonymous):

3x+5y<10 rearrange to get 5y<10-3x and again to \[y < 2 - \frac{ 3 }{ 5 } x\] Please rewrite the second equation, I can't see what inequality it is between the 4y and 12.

OpenStudy (anonymous):

this is what I have s far x<10/3-5y/3 and x<- 4y+12

OpenStudy (anonymous):

ok so for the second equation 4y<12-x -> \[y < 3 - \frac{ x }{ 4 }\] If you pretend that those inequality signs are equals signs, are you able to draw those graphs?

OpenStudy (anonymous):

nope I am having trouble with graphing

OpenStudy (anonymous):

For each equation pick a few x terms (say 0, 1 and 2 to make it easy) and see what y value you get. Then plot the 3 points and draw a line through them.

OpenStudy (anonymous):

ok I figured out the fisrt line

OpenStudy (anonymous):

the sectond line is supposed to be dashed right?

OpenStudy (anonymous):

I'm not sure what you've graphed there but I don't think it's any of the equations you had. Hint: when the equation of the line is y=mx+c, c is the y intercept and m is the gradient (slope). One should cross the y axis at 2 and the other at 3.

OpenStudy (anonymous):

here are the posible graphs

OpenStudy (anonymous):